OpenStudy (zyberg):

A room gives out 87 MJ heat every day. A company wants to make an electrical heater to maintain a constant temperature in the room. How long of a nichrome 1mm diameter wire would heating element need?

OpenStudy (zyberg):

I am completely at loss here. I think that at some point the formula \[R = \frac{length * ro}{diameter}\] (ro as in a greek letter) would be useful, but I have no idea how to convert heat into anything useful... @ganeshie8

OpenStudy (irishboy123):

that'll give you the resistance for a certain length and x-sect area you need more info - eg voltage

OpenStudy (irishboy123):

link the actual question if you can. it's based on the same idea as the others you posted earlier but info is missing :-(

OpenStudy (zyberg):

Well, info isn't missing. That's why I posted this here.

OpenStudy (zyberg):

I think that, perhaps, we could figure something out of a word "constant temperature", but I am not sure. Also, it is possible that the answer should be expressed in letters (although I doubt that).

OpenStudy (zyberg):

Just googled, seems that ro is between 1 and 1.5 * 10^-6. We can use constants.

OpenStudy (zyberg):

So, the only thing that is missing - how to take out resistance out of heat?

OpenStudy (irishboy123):

in letters if you are using a fixed voltage \(V\), the power is \(P = \dfrac{V^2}{R}\) -- that's what they mean, the power loss is heat. or \(P = IV = I^2 R\) So with \(R = \dfrac{\rho L }{A}\) so \(\Large P = \dfrac{V^2}{\dfrac{\rho L }{A}} = 87*10^6\) with \(\Large A = \dfrac{\pi (0.001)^2}{4}\) you an find the resistivity online somewhere so....play with that and L is the only unknown provided you have a V

OpenStudy (irishboy123):

Oh crap. Howler alert!! "A room gives out 87 MJ heat **every day**." So power P is actually \(P = \dfrac{87.10^6}{24*60*60} ~ J/s\) So the power the heater must supply to compensate is \(\large P = \dfrac{V^2}{\dfrac{\rho L }{A}} \approx 1007\) and solve for L but you still have an unknown in there. you need to know the voltage or current